System Status
All Systems Operational
Domain Services Operational
Core gTLDs   ? Operational
Core ccTLDs   ? Operational
Other gTLDs   ? Operational
Other ccTLDs   ? Operational
Premium Names   ? Operational
WHOIS   ? Operational
Test Environment Domain Services   Operational
APIs Operational
OpenSRS API   ? Operational
OpenHRS API   Operational
Email API   ? Operational
Test Environment API   Operational
Hosted Email Operational
Cluster A   Operational
Cluster B   Operational
Webmail   Operational
Control Panels Operational
Reseller Control Panel   Operational
Classic RWI   ? Operational
End User Control Panel   Operational
Storefront   Operational
Payment Gateway   ? Operational
Test Environment Control Panels   Operational
DNS Operational
SystemDNS   Operational
Domain Forwarding   Operational
SSL   ? Operational
Reseller Support   ? Operational
GoMobi   Operational
System Messaging   ? Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Scheduled Maintenance
OpenSRS Platform Sep 27, 04:00-07:00 UTC
We are performing proactive firmware upgrades to our load balancing infrastructure in order to stay ahead of security vulnerabilities and other bugs. During this maintenance window, domain management services, including registration, transfer, and trade functionality, may be unavailable, or their performance may be degraded. Also, registrants may be unable to access their Consent Management Page URL or complete Whois verification. Downtime is not expected to exceed 15 minutes, and service interruptions are unlikely.

If you have any questions or concerns, please don't hesitate to contact us at help@opensrs.com.
Posted on Sep 14, 11:33 UTC
OpenSRS Platform Oct 3, 02:00-03:00 UTC
We are proactively making updates to our infrastructure in order to improve stablilty within our systems. During this maintenance window, domain management services, including registration, renewal, transfer, and update functionality, may be unavailable, or their performance may be degraded. Also, users may experience errors when accessing their Reseller Control Panel and when sending API calls.

We apologize for the inconvenience, and we appreciate your patience. If you have any questions or concerns, please don't hesitate to contact us at help@opensrs.com.
Posted on Sep 20, 19:43 UTC
Nominet Registry Oct 4, 05:00-07:00 UTC
During the maintenance all domains will continue to resolve, but all other services may be unavailable.

For a list of affected TLD's please look the registry up in the following spreadsheet:
http://opensrs.help/chart
Posted on Sep 11, 08:47 UTC
Nominet Registry Oct 9, 05:00-07:00 UTC
During the maintenance all domains will continue to resolve, but all other services may be unavailable.

For a list of affected TLD's please look the registry up in the following spreadsheet:
http://opensrs.help/chart
Posted on Sep 11, 08:48 UTC
.CH + .LI Registry Oct 11, 05:00-05:30 UTC
During the maintenance all domains will continue to resolve, but all other services may be unavailable.
Posted on Jul 23, 13:04 UTC
Verisign Registry Oct 18, 21:00-23:00 UTC
During the maintenance all domains will continue to resolve, but all other services may be unavailable.

For a list of affected TLD's please look the registry up in the following spreadsheet:
http://opensrs.help/chart
Posted on Sep 20, 07:14 UTC
.FR Registry Nov 19, 07:00-11:00 UTC
During the maintenance all domains will continue to resolve, but all other services may be unavailable.
Posted on Sep 6, 20:52 UTC
Past Incidents
Sep 26, 2018

No incidents reported today.

Sep 25, 2018
Completed - The scheduled maintenance has been completed.
Sep 25, 21:15 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Sep 25, 19:00 UTC
Scheduled - During the maintenance all domains will continue to resolve, but all other services may be unavailable.
Sep 20, 07:34 UTC
Completed - The scheduled maintenance has been completed.
Sep 25, 07:00 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Sep 25, 03:00 UTC
Scheduled - During the maintenance all domains will continue to resolve, but all other services may be unavailable.

For a list of affected TLD's please look the registry up in the following spreadsheet:
http://opensrs.help/chart
Sep 20, 08:00 UTC
Sep 24, 2018
Completed - The scheduled maintenance has been completed.
Sep 24, 17:30 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Sep 24, 17:01 UTC
Scheduled - During the maintenance all domains will continue to resolve, but all other services may be unavailable.

For a list of affected TLD's please look the registry up in the following spreadsheet:
http://opensrs.help/chart
Sep 14, 19:25 UTC
Sep 23, 2018
Resolved - The Verisign registry underwent production maintenance for the .COM and .NET TLDs tonight from approximately 9pm ET to 9:45pm et. All .COM and .NET domains would have resolved fine, but all other domain management services for those TLDs may have been unavailable during that period but full functionality has now been restored.
Sep 23, 02:41 UTC
Sep 22, 2018
Resolved - The issues we were experiencing connecting to the ARI registry proxy have now been resolved. They were due to an issue with a certificate installation but the full management capabilities of ARI managed TLD domains for both SRS and HRS resellers has now been restored.
Sep 22, 17:12 UTC
Investigating - We are currently experiencing issues connecting to the ARI registry proxy. This will impact the placement of new domain orders, and the management of existing ARI TLD domains. You can confirm the TLDs managed by ARI via our TLD Reference Chart.

https://docs.google.com/spreadsheets/d/13t4l-kO3qAio4RCF3j1lF0X2AxaIr_G5kFIqIF3LAZU/edit#gid=1203932299
Sep 22, 02:51 UTC
Sep 21, 2018
Resolved - -
We are happy to report that our connection to Comodo has been restored, and you may place any pending orders.
Sep 21, 18:56 UTC
Identified - -
We are working on a fix for this and will update again once we receive additional information.
Sep 21, 13:33 UTC
Investigating - We are currently experiencing issues connecting to the ceritifcate authority Comodo. This would impact the placement of new orders and any management of any existing pending orders. We have reached out to Comodo and we expect to have this resolved shortly. This only impacts the ordering of Comodo SSL certificates. All other certificates still function fine. We will provide an update as soon as we have more information.
Sep 21, 00:37 UTC
Sep 20, 2018
Resolved - We encountered a brief disk failure at 5:22pm ET. It was idenitifed almost immediately and restored, however, during the recovery process from then until approximately 5:45pm ET, users of the Reseller Control Panel may have experienced intermittent issues logging in, and navigating through and loading certain pages and elements. These issues are now resolved with full functionality restored.
Sep 20, 21:56 UTC
Resolved - -
We are happy to report that we are no longer seeing any reports or alerts for our test enviornment. All services are operational.
Sep 20, 19:36 UTC
Monitoring - -
Our operations team has identified the cause and made the appropriate corrections. The result was from a recent promo which was rolled back.
We are standing in monitoring until further notice.
Sep 20, 19:22 UTC
Investigating - -
We are experiencing issues with our testing environment. You may notice errors with domain services and connections to the testing environment.

Our operations team is currently investigating the issue and we will post additional information once received.
Sep 20, 18:14 UTC
Completed - The scheduled maintenance has been completed.
Sep 20, 05:30 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Sep 20, 05:00 UTC
Scheduled - During the maintenance all domains will continue to resolve, but all other services may be unavailable.
Jul 23, 13:02 UTC
Sep 19, 2018
Completed - The scheduled maintenance has been completed.
Sep 19, 20:00 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Sep 19, 17:00 UTC
Scheduled - During the maintenance all domains will continue to resolve, but all other services may be unavailable.

For a list of affected TLD's please look the registry up in the following spreadsheet:
http://opensrs.help/chart
Sep 11, 08:55 UTC
Resolved - -
We are happy to report the listed TLDs are now operational and transfer orders should now be processing as desired.
Please proceed with processing any previous pending orders.
Sep 19, 19:03 UTC
Monitoring - -
We have confirmed the issue and have implemented a fix for this. We will be monitoring the status for the listed TLDs.
We will update to confirm completion once additional data is received.
Sep 19, 17:56 UTC
Update - -
Our team continues to work on this issue. You may receive an error message:

Order ######### () not completed in the registry: Transfer failed due to internal server problems.

While placing a transfer order for the listed TLDs. More information will be provided when we receive more data.
Sep 19, 14:49 UTC
Investigating - -
We are currently experiencing intermittent issues with the transfer in of domains under the .AU, .DE, .CH, & .SE TLDs. We are investigating the matter and we will provide more information as soon as possible.
Sep 18, 23:41 UTC
Sep 18, 2018
Completed - The scheduled maintenance has been completed.
Sep 18, 20:01 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Sep 18, 17:00 UTC
Scheduled - During the maintenance all domains will continue to resolve, but all other services may be unavailable.

For a list of affected TLD's please look the registry up in the following spreadsheet:
http://opensrs.help/chart
Sep 11, 08:54 UTC
Completed - The scheduled maintenance has been completed.
Sep 18, 15:30 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Sep 18, 14:00 UTC
Scheduled - During the maintenance all domains will continue to resolve, but all other services may be unavailable.
Sep 10, 08:03 UTC
Sep 17, 2018
Resolved - This incident has been resolved.
Sep 17, 09:17 UTC
Monitoring - Our operations team has resolved the issues which impacted domain registrations, contact updates and whois lookups between 05:17 UTC and 08:16 UTC. We will continue to monitor for any further issues.
Sep 17, 08:48 UTC
Update - Services are now recovering. There is still an intermittent issue with our whois service which is being addressed by our operations team.
Sep 17, 08:09 UTC
Update - Our operations team continues work to address the issue affecting registrations, contact information management and whois lookups. There are no new updates at this time.
Sep 17, 07:32 UTC
Update - There are no additional updates at this time. Our operations team is still working towards a resolution.
Sep 17, 06:58 UTC
Identified - Complications are currently being dealt with by Operations and issues with processing domains registrations, contact updates, as well as Whois lookups can be expected until resolved. This issue began at 05:17 UTC and is still being worked on.
Sep 17, 06:25 UTC
Investigating - We are investigating reports of API registrations, contact updates and Whois checks potentially failing. DNS resolution is not affected at this time. Our operations team has been engaged.
Sep 17, 05:54 UTC
Sep 16, 2018
Completed - The scheduled maintenance has been completed.
Sep 16, 10:00 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Sep 16, 06:00 UTC
Scheduled - During the maintenance all domains will continue to resolve, but all other services may be unavailable.
Sep 14, 19:21 UTC
Sep 15, 2018
Resolved - This incident has been resolved.
Sep 15, 13:00 UTC
Update - -
Services have been restored and systems are operational, however, we will continue to monitor to ensure there are no further issues.
Sep 13, 20:22 UTC
Monitoring - -
The DDOS stopped at 12:05 UTC and services are now recovering. Some API issues are still being addressed, but the majority of the issues have been resolved. We will continue to monitor for any further issues.
Sep 13, 13:01 UTC
Update - -
We have identified the root cause of the outage as a DDOS and are working with our upstream network providers to resolve. This issue has the highest level of attention internally and we will provide additional updates as they become available
Sep 13, 11:42 UTC
Identified - -
The issue affecting OpenSRS control panels and API is still being mitigated by our operations team. There are no additional updates at this time.
Sep 13, 11:18 UTC
Investigating - -
At this time the Reseller Control Panel (RCP/UCP), Classic Reseller Interface (RWI), and Admin Web Interface (AWI) may be unavailable. Domain orders via the API are also affected.

This is currently being investigated by our operations team.
Sep 13, 09:37 UTC
Resolved - We experienced intermittent issues with IMAP and POP3 Connectivity on our Cluster A from approximately 7:51pm ET to 8:03pm ET due to a disk issue. Some users may have experienced difficulties authenticating via IMAP or POP3 as a result, however, the issues were quickly identified and resolved with services now back online fully restored.
Sep 15, 00:36 UTC
Sep 14, 2018
Resolved - There have been no additional issues since the last update. The delays have cleared up and the issue has been resolved.
Sep 14, 13:54 UTC
Monitoring - Our operations team has adjusted filtering for outbound mail on both clusters. Monitoring will continue until delays have fully subsided.
Sep 14, 13:26 UTC
Investigating - We are investigating a potential issue with sending mail through both cluster A and B, for Hostedemail. Users may experience slowness or an error when sending emails. Our operations team has been engaged.
Sep 14, 12:39 UTC
Completed - The scheduled maintenance has been completed.
Sep 14, 06:00 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Sep 14, 03:00 UTC
Scheduled - During the maintenance all domains will continue to resolve, but all other services may be unavailable.

For a list of affected TLD's please look the registry up in the following spreadsheet:
http://opensrs.help/chart
Sep 11, 08:52 UTC
Sep 13, 2018
Resolved - The intermittent issues with SMTP outbound email on Cluster B are now resolved with full regular email service restored to the cluster. Our Operations team confirmed the issues were due to a pool of users on the cluster engaged in abusive email behavior, and they have now all been blocked / suspended.
Sep 13, 02:14 UTC
Update - We are still experiencing intermittent issues with SMTP outbound email on Cluster B, as our Operations team continues to troubleshoot the issue. However, at this point the issues should be isolated to just intermittent slowness when sending messages, so we hope to have this completely resolved soon. We will update you as soon as we have more information to provide, and we appreciate your patience.
Sep 13, 01:06 UTC
Investigating - We have a handful of reports that SMTP connectivity for cluster B may be experiencing intermittent downtimes.

More updates will become available soon
Sep 12, 19:29 UTC